OpenStudy (banana):

true or false a^3*a^5=a^15

OpenStudy (anonymous):

false \[a^m*a^n=a^{m+n}\]

OpenStudy (banana):

(b^2)^4=b^6

OpenStudy (anonymous):

no \[(b^m)^n= b^{m*n}\]

OpenStudy (amistre64):

my vote is....false

OpenStudy (banana):

(x+y)^-2=x^-2+y^-2

OpenStudy (anonymous):

also false

OpenStudy (banana):

simply 4c/12c^5

OpenStudy (banana):

3x-y=2 16x-5y=11 parallel or coinciding